At the beginning of a walk, Roberto and Juan are 7.7 miles apart. If they leave at the same time Roberto meets Juan in 11 hours. If they walk towards each other they meet in 1 hour. What are their average speeds?
I need to know how to set up these equations. I don't even know where to begin.

Answer by mananth(16946)

at the same time Roberto meets Juan in 11 hours. If they walk towards each other they meet in 1 hour. What are their average speeds?

Roberto speed = x mph
Juan speed = y mph
let x>y
d= 7.7
if they go away then the catch up distance = 7.7
the catch up speed = x-y
t= d/r
t=11
11= 7.7/(x-y)
11x-11y=7.7
/11
x-y = 0.7..............1
.....
if they move towards
then speed = x+y
t=1
d= 7.7
t=d/r
1=7.7/(x+y)
x+y = 7.7 ------------2
add (1) & (2)
2x=0.7 + 7.7
2x=8.4
/2
x=4.2
Roberto speed = 4.2 mph
from (2) x+y =7.7
4.2+y=7.7
y=7.7-4.2
y=3.5
Juan speed = 3.5 mph
m.ananth@hotmail.ca